Please help me to to solve this quadratic equation. Solve each equation in part A using factoring. Solve the equation in part b by completing the square
Part A. 8x^2 +2x - 3 = 0
Part B. 8x^2 +2x - 2 = 0

Answer by ewatrrr(24785)

F First terms
O Outside terms + I Inside terms
L Last terms
......
8x^2 +2x - 3 = 0
(4x ± _)(2x ± _) = 0
(4x+ 3)(2x -1)= 0, x= -3/4, 1/2
..........
8x^2 +2x - 2 = 0 Format ax^2 + bx + c
b/-2a = -1/8
8(x - (- 1/8))^2 - 8( 1/64) - 2 = 0
8(x + 1/8)^2 - 1/8 - 2 = 0
8(x + 1/8)^2 - 9/8 = 0
(x+1/8)^2 = 9/64
......
x = -1/8 ± 3/8
x = -4/8 = -1/2
0r
x = 2/8 = 1/4
........
Check by factoring
8x^2 +2x - 2 = 0
(4x-1)(2x+1)= 0